OpenStudy (anonymous):

Use a linear approximation to estimate the number √9.05 .

OpenStudy (anonymous):

you can probably use \[f(x)=\sqrt{9+x}\] and linearize that one then use \(x=.05\)

OpenStudy (anonymous):

\[f(x)=\sqrt{9+x}\] at \((0,3)\) \[f'(x)=\frac{1}{2\sqrt{9+x}}\]\[f'(0)=\frac{1}{6}\]

OpenStudy (anonymous):

your lineariziation is \[y=\frac{1}{6}x+3\]

OpenStudy (anonymous):

then finish by replacing \(x\) by \(.05\)